OpenStudy (anonymous):

The diameter of a construction pipe is 4ft . If the pipe is 17ft long, what is its volume?

OpenStudy (anonymous):

pi*2^2*17

OpenStudy (anonymous):

so 213.52 ft^3

OpenStudy (nurali):

r = 4/2 = 2 ft V=pi*r^2*l V=pi*(2)^2*17 V=68*pi cubic feet